OpenStudy (anonymous):

How can x + 8 = −x + 7 be set up as a system of equations?

OpenStudy (anonymous):

x+8=? -x+7=? Obsiously the first and second problem are the same, so you can say: -x+7=x+8 or (-8)-(-x)= x-(-7). Does this help you? http://www.purplemath.com/modules/systlin4.htm hopefully this site will help